ckitookthw0ck2poland:

Solve the proportion. Where necessary, round to the nearest hundredth. 8/2 = x+3/4 A. 13 B. 1 C. 7 D. 16

Laylalyssa:

4= x+3/4 x+3/4=4 then subtract 3/4 from both sides and then u get x= 13/4

ckitookthw0ck2poland:

THANK U SOOO MUCH LAYLAAAAAAAAAAAAA

Laylalyssa:

you're welcomeee <3